# Decoupled structure-preserving discretization of incompressible MHD equations with general boundary conditions

### v1

@@ Abstract
In the framework of a mixed finite element method, a structure-preserving formulation for incompressible MHD equations with general boundary conditions is proposed. A leapfrog-type temporal scheme fully decouples the fluid part from the Maxwell part by means of staggered discrete time sequences and, in doing so, partially linearizes the system. Conservation and dissipation properties of the formulation before and after the decoupling are analyzed. We demonstrate optimal spatial and second-order temporal error convergence and conservation and dissipation properties of the proposed method using manufactured solutions, and apply it to the benchmark Orszag-Tang and lid-driven cavity test cases.
@@
MHD , mixed finite element method , structure-preserving discretization , general boundary conditions
@@
[1]organization=School of Mathematics and Computing Science, Guilin University of Electronic Technology, city=Guilin, country=China
@@ 1 Introduction
The incompressible MHD equations, (1), are a coupled system of the incompressible Navier-Stokes equations and the Maxwell’s equations. The two evolution equations, (1a) and (1c), stand for the conservation of momentum and the Faraday’s law of induction, respectively. Ohm’s law and Ampere’s law are (1d) and (1e), respectively. (1f) sets up a constitutive relation between $\bm{B}$ and $\bm{H}$. The divergence-free constraint on $\bm{u}$, (1b), represents mass conservation of the constant mass density flow. It is seen that, as $\bm{B}^{0}:=\bm{B}(\bm{x},0)$ is divergence free, (1c) ensures
@@
which is Gauss’s law for magnetism. And, from (1e), we know $\nabla\cdot\bm{j}=0$ which stands for conservation of charge. The coupling between the flow field and the electromagnetic field is through the Lorentz force, $\bm{j}\times\bm{B}$ in (1a), and the electromotive force, $\bm{u}\times\bm{B}$ in (1d).
@@
Note that we do not introduce new notations for the scaled objects and, after this point, they always refer to the scaled dimensionless ones.
Now (1) can be written in a dimensionless form, i.e.,
@@
have been written into (2d), which is the magnetic induction (or Helmholtz) equation. [^1] Note that, after non-dimensionalization, we get $\bm{B}=\bm{H}$, i.e. (3d). This is the case in vector calculus. In differential forms, they appear as a pair of differently oriented dual forms connected by the Hodge $\star$ operator, i.e., $\bm{B}=\star\bm{H}$. The three dimensionless parameters are the fluid Reynolds number $\mathrm{R}_{f}=\dfrac{\rho UL}{\tilde{\mu}}=\dfrac{UL}{\nu}$ (with $\nu=\dfrac{\tilde{\mu}}{\rho}$ being the kinematic viscosity), the magnetic Reynolds number $\mathrm{R}_{m}=\mu\sigma UL$, and the coupling number $\mathsf{c}=\mathrm{A}_{l}^{-2}$ ($\mathrm{A}_{l}=\dfrac{U\sqrt{\rho\mu}}{B}=\dfrac{U}{U_{\mathrm{A}}}$ is the Alfvén number and $U_{\mathrm{A}}=\dfrac{B}{\sqrt{\rho\mu}}$ is the Alfvén speed). Sometimes, the coupling term in (2a), $\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}$, is written as $\mathsf{s}\left(\bm{E}+\bm{u}\times\bm{H}\right)\times\bm{H}$ where the coupling number becomes $\mathsf{s}=\mathsf{c}\mathrm{R}_{m}=\dfrac{\sigma LB^{2}}{\rho U}$. For example, see [3, (2.2a)].
@@
    (4) {$&P=\widehat{P}\quad&&\text{on }\Gamma_{\widehat{P}}\times\left(0,T\right]\\ &\bm{u}\cdot\bm{n}=\widehat{u}\quad&&\text{on }\Gamma_{\widehat{u}}\times\left(0,T\right]\end{aligned}\right.\quad\text{and}\quad\left\{\begin{aligned} &\bm{u}\times\bm{n}=\widehat{\bm{u}}\quad&&\text{on }\Gamma_{\widehat{\bm{u}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{\omega}\times\bm{n}\right)=\widehat{\bm{\omega}}\quad&&\text{on }\Gamma_{\widehat{\bm{\omega}}}\times\left(0,T\right]\end{aligned}\right.,$
    where $\bm{n}$ is the outward unit normal vector, $\left\{\Gamma_{\widehat{P}},\Gamma_{\widehat{u}}\right\}$ and $\left\{\Gamma_{\widehat{\bm{u}}},\Gamma_{\widehat{\bm{\omega}}}\right\}$ are two partitions of the boundary, such that $\overline{\Gamma}_{\widehat{P}}\cup\overline{\Gamma}_{\widehat{u}}=\partial\Omega$, $\Gamma_{\widehat{P}}\cap\Gamma_{\widehat{u}}=\emptyset$ and $\overline{\Gamma}_{\widehat{\bm{u}}}\cup\overline{\Gamma}_{\widehat{\bm{\omega}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{u}}}\cap\Gamma_{\widehat{\bm{\omega}}}=\emptyset$.
@@
    on $\partial\Omega\times\left(0,T\right]$. Also see [4, Section 2.3].
@@
For example, the boundary conditions considered in the work of Hu et al., see (7) of [5], are a particular case of this general configuration, i.e. the case when $\Gamma_{\widehat{u}}=\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset$, $\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$.
@@
It is generally acknowledged that structure-preservation, especially the preservation (either strongly of weakly) of $\nabla\cdot\bm{B}=0$, is an essential property of numerical methods for MHD [6]. Many of these structure-preserving methods can be classified into, for example, potential-based methods [7], divergence-cleaning methods [8, 9], constrained transport methods [10, 11] and Lagrange multiplier (or augmented) methods [12, 13, 14]. In the current paper neither of those techniques will be applied, but it will be shown that the absence of magnetic monopoles is weakly enforced through the decoupled variational formulation. Recently, methods that aim at preserving multiple structures in the mixed finite element setting become increasingly popular. See, for example, the work of Hu et al. for incompressible MHD that preserves cross- and magnetic-helicity, energy and Gauss law of magnetism [5], the work of Gawlik and Gay-Balmaz for incompressible MHD of a variable fluid density that preserves energy, cross-helicity (when the fluid density is constant) and magnetic-helicity, mass, total squared density, pointwise incompressibility, and Gauss’s law for magnetism [15], the work of Laakmann et al. [16] and references therein. These methods preserve several physical quantities of interest, but usually are less computationally efficient because of their large (nonlinear) discrete systems to be solved. A compromise is to use a decoupled scheme which breaks the discrete system into several smaller ones, for example see [3, 12, 13, 17, 18]. These methods in the mixed finite element setting commonly deal with a particular configuration of boundary conditions.
@@
In this paper, we introduce a structure-preserving mixed finite element discretization for the incompressible MHD under a general boundary condition configuration. The method is decoupled in the sense that the two evolution equations are temporally discretized at different time sequences such that they can be solved individually.
@@
The outline of the rest of the paper is as follows. Relevant Hilbert spaces and the de Rham Hilbert complex are introduced in Section 2 followed by the proposed semi-discrete formulation and its conservation properties in Section 3. The temporal discretization that leads to the decoupled scheme is explained in Section 4. Numerical tests are presented in Section 5. Finally, in Section 6 conclusions are drawn.
@@ 2 Relevant Hilbert spaces and the de Rham complex
The function spaces used in this work are similar to those used in [19] where only the two-dimensional setting is discussed. For the sake of completeness, we give a brief introduction to the three-dimensional spaces as a complement to [19, Section 2.1]. We first introduce the relevant spaces and the de Rham complex in the infinite dimensional setting, and then introduce their finite dimensional counterparts. For an extensive introduction of function spaces, see [20] or any other textbook on functional analysis.
2.1 The infinite dimensional setting
@@
They form the well known de Rham Hilbert complex [21, 22],
(7) $\mathbb{R}\hookrightarrow H^{1}\left(\Omega\right)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}H(\mathrm{curl};\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}H(\mathrm{div};\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}L^{2}(\Omega)\to 0.$
@@
Note that, at the continuous level, although $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ are different, see (8), trace spaces $\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right)$ and $\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right)$ refer to the same tangential space.
2.2 The finite dimensional setting
Assume that we are given finite dimensional spaces $G(\Omega)\subset H^{1}(\Omega)$, $C(\Omega)\subset H(\mathrm{curl};\Omega)$, $D(\Omega)\subset H(\mathrm{div};\Omega)$, $S(\Omega)\subset L^{2}(\Omega)$, which form a discretization of the continuous de Rham Hilbert complex,
(9) $\mathbb{R}\hookrightarrow G(\Omega)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}C(\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}D(\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}S(\Omega)\to 0.$
In particular, this means that if the complex in (7) is exact, then so is the one in (9). We will also consider the following subspaces with boundary conditions, where $\Gamma\subset\partial\Omega$,
@@ 2.2 The finite dimensional setting
Moreover, the finite dimensional trace spaces will be denoted by $\mathcal{T}G(\Omega,\Gamma)$, $\mathcal{T}C(\Omega,\Gamma)$, $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$, $\mathcal{T}D(\Omega,\Gamma)$. Unlike their continuous counterparts, finite dimensional trace spaces $\mathcal{T}C(\Omega,\Gamma)$ and $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$ usually are different.
@@
which is skew-symmetric with respect to any two of the three entries. Note that, in the infinite dimensional setting, the $L^{2}$-boundedness of $\left\langle{\bm{\alpha}\times\bm{\beta}},{\bm{\gamma}}\right\rangle_{\Omega}$ is not implied even if $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are $L^{2}$ vector fields. However, in the finite dimensional setting where the spaces $C(\Omega)$ and $D(\Omega)$ consist of piecewise-polynomial finite element vector fields, it is in fact $L^{2}$-bounded.
@@ 3 Spatially-discrete formulation and its conservation properties
In this section, we present a spatially-discrete weak formulation of (2) and analyze its conservation properties. As a remark, for neatness, we only discuss the three-dimensional formulation and its temporal discretization in this paper. Obtaining their two-dimensional versions is straightforward.
@@ 3.1 Spatially-discrete formulation
To increase the notational clarity, we omit the part for the temporal domain. The proposed semi-discrete weak formulation of (2) in $\mathbb{R}^{3}$ is written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, seek $\left(\bm{u}_{h},\bm{\omega}_{h},P_{h},\bm{H}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}_{\perp}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
@@
Note that this assumes that the supplied boundary conditions are exactly representable as members of the chosen finite dimensional spaces; more generally, one can always project the boundary conditions into the appropriate finite dimensional trace spaces. See (4) and (5) for the general boundary setting.

### v2

@@ Abstract
In the framework of a mixed finite element method, a structure-preserving formulation for incompressible magnetohydrodynamic (MHD) equations with general boundary conditions is proposed. A leapfrog-type temporal scheme fully decouples the fluid part from the Maxwell part by means of staggered discrete time sequences and, in doing so, partially linearizes the system. Conservation and dissipation properties of the formulation before and after the decoupling are analyzed. We demonstrate optimal spatial and second-order temporal accuracy, as well as conservation and dissipation properties, of the proposed method using manufactured solutions, and apply it to the benchmark Orszag-Tang and lid-driven cavity cases.
@@
MHD , structure-preserving discretization , decoupled discretization , mixed finite element method , general boundary conditions
@@
[1]organization=Guilin University of Electronic Technology, School of Mathematics and Computing Science, city=Guilin, country=China
@@ 1 Introduction
The incompressible MHD equations, (1), are a coupled system of the incompressible Navier-Stokes equations and Maxwell’s equations. The two evolution equations, (1a) and (1c), stand for the conservation of momentum and Faraday’s law of induction, respectively. Ohm’s law and Ampere’s law are (1d) and (1e), respectively. (1f) sets up a constitutive relation between $\bm{B}$ and $\bm{H}$. The divergence-free constraint on $\bm{u}$, (1b), represents mass conservation of the constant mass density flow. It is seen that, as $\bm{B}^{0}:=\bm{B}(\bm{x},0)$ is divergence-free, (1c) ensures
@@
which is Gauss’s law for magnetism. And, from (1e), we know $\nabla\cdot\bm{j}=0$ which stands for conservation of charge. The coupling between the flow field and the electromagnetic field is achieved through the terms $\bm{j}\times\bm{B}$ in (1a) and $\bm{u}\times\bm{B}$ in (1d).
@@
Note that we do not introduce new notation for the scaled objects, and, after this point, they always refer to the scaled dimensionless ones.
Now (1) can be written in a dimensionless form,
@@
have been written into (2d), which is the magnetic induction (or Helmholtz) equation. [^1] Note that, after non-dimensionalization, we get $\bm{B}=\bm{H}$, i.e. (3d). This is the case in vector calculus. In differential forms, they appear as a pair of differently oriented dual forms connected by the Hodge $\star$ operator, i.e. $\bm{B}=\star\bm{H}$. The three dimensionless parameters are the fluid Reynolds number $\mathrm{R}_{f}=\dfrac{\rho UL}{\tilde{\mu}}=\dfrac{UL}{\nu}$ (with $\nu=\dfrac{\tilde{\mu}}{\rho}$ being the kinematic viscosity), the magnetic Reynolds number $\mathrm{R}_{m}=\mu\sigma UL$, and the coupling number $\mathsf{c}=\mathrm{A}_{l}^{-2}$ ($\mathrm{A}_{l}=\dfrac{U\sqrt{\rho\mu}}{B}=\dfrac{U}{U_{\mathrm{A}}}$ is the Alfvén number and $U_{\mathrm{A}}=\dfrac{B}{\sqrt{\rho\mu}}$ is the Alfvén speed). Sometimes, the coupling term in (2a), $\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}$, is written as $\mathsf{s}\left(\bm{E}+\bm{u}\times\bm{H}\right)\times\bm{H}$ where the coupling number becomes $\mathsf{s}=\mathsf{c}\mathrm{R}_{m}=\dfrac{\sigma LB^{2}}{\rho U}$. For example, see [3, (2.2a)]. One can see that, from (3),
$\partial_{t}\bm{H}=\partial_{t}\bm{B}=-\nabla\times\bm{E},$
and due to this, we will have that
(4) $\partial_{t}\nabla\cdot\bm{H}=-\nabla\cdot\nabla\times\bm{E}=0,$
which leads to Gauss’s law for magnetism if $\bm{H}^{0}$ is divergence-free.
@@
    (5) {$&P=\widehat{P}\quad&&\text{on }\Gamma_{\widehat{P}}\times\left(0,T\right]\\ &\bm{u}\cdot\bm{n}=\hat{u}\quad&&\text{on }\Gamma_{\hat{u}}\times\left(0,T\right]\end{aligned}\right.\quad\text{and}\quad\left\{\begin{aligned} &\bm{u}\times\bm{n}=\widehat{\bm{u}}\quad&&\text{on }\Gamma_{\widehat{\bm{u}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{\omega}\times\bm{n}\right)=\widehat{\bm{\omega}}\quad&&\text{on }\Gamma_{\widehat{\bm{\omega}}}\times\left(0,T\right]\end{aligned}\right.,$
    where $\bm{n}$ is the outward unit normal vector, $\left\{\Gamma_{\widehat{P}},\Gamma_{\hat{u}}\right\}$ and $\left\{\Gamma_{\widehat{\bm{u}}},\Gamma_{\widehat{\bm{\omega}}}\right\}$ are two partitions of the boundary, such that $\overline{\Gamma}_{\widehat{P}}\cup\overline{\Gamma}_{\hat{u}}=\partial\Omega$, $\Gamma_{\widehat{P}}\cap\Gamma_{\hat{u}}=\emptyset$ and $\overline{\Gamma}_{\widehat{\bm{u}}}\cup\overline{\Gamma}_{\widehat{\bm{\omega}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{u}}}\cap\Gamma_{\widehat{\bm{\omega}}}=\emptyset$.
@@
    on $\partial\Omega\times\left(0,T\right]$. Also, see [4, Section 2.3].
@@
For example, the boundary conditions considered in the work of Hu et al., see (7) of [5], are a particular case of this general configuration, i.e. the case when $\Gamma_{\hat{u}}=\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset$, $\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$.
@@
It is generally acknowledged that structure-preservation, especially the preservation (either strongly or weakly) of $\nabla\cdot\bm{B}=0$, is an essential property of numerical methods for MHD [6]. Many of these structure-preserving methods can be classified into, for example, potential-based methods [7], divergence-cleaning methods [8, 9], constrained transport methods [10, 11] and Lagrange multiplier (or augmented) methods [12, 13, 14]. In the current paper, none of those techniques will be applied, but it will be shown that the absence of magnetic monopoles is weakly enforced through the decoupled variational formulation.
Recently, methods that aim at preserving multiple structures in the mixed finite element setting have become increasingly popular. See, for example, the work of Hu et al. for incompressible MHD that preserves cross- and magnetic-helicity, energy, and Gauss law of magnetism [5], the work of Gawlik and Gay-Balmaz for incompressible MHD of a variable fluid density that preserves energy, cross-helicity (when the fluid density is constant) and magnetic-helicity, mass, total squared density, pointwise incompressibility, and Gauss’s law for magnetism [15], the work of Laakmann et al. [16] and references therein. These methods preserve several physical quantities of interest, but are usually less computationally efficient because of their large (nonlinear) discrete systems to be solved.
A compromise is to use a decoupled scheme that breaks the discrete system into several smaller ones. For example, Zhang and Su [3] propose a three-step decoupled scheme by using a first-order semi-implicit Euler scheme with a first-order stabilizer and some carefully designed implicit-explicit treatments for the coupling terms. In [13], a four-step decoupled scheme of second-order temporal accuracy is studied. It employs the mixed finite element method under a “zero-energy-contribution” idea and a projection method. Ma and Huang introduced a vector penalty-projection method, a five-step decoupled scheme that has a mixed finite element spatial discretization and a first-order backward Euler temporal discretization in [17]. The work [18] reports another five-step decoupled scheme which combines the idea of an auxiliary variable, pressure correction projection method, and some other implicit-explicit techniques for the thermally coupled MHD equations with variable density. For a more extended literature study on decoupled schemes in a mixed finite element setting for MHD equations, we refer to [17]. These methods commonly deal with a particular configuration of boundary conditions.
@@
In this paper, we introduce a mixed finite element structure-preserving decoupled discretization of a second-order temporal accuracy for the incompressible MHD equations with general boundary conditions. This method decouples each full iteration into two steps by means of an implicit mid-point scheme that is applied to the two evolution equations at staggered temporal sequences such that they can be solved individually in a leapfrog approach.
@@
The outline of the rest of the paper is as follows. Relevant Hilbert spaces and the de Rham Hilbert complex are introduced in Section 2 followed by the proposed semi-discrete formulation and its conservation properties in Section 3. The temporal discretization that leads to the decoupled scheme is explained in Section 4. Numerical tests are presented in Section 5. Finally, in Section 6, conclusions are drawn.
@@ 2 Relevant Hilbert spaces and the de Rham complex
The function spaces used in this work are similar to those used in [19] where only the two-dimensional setting is discussed. For the sake of completeness, we give a brief introduction to the three-dimensional spaces as a complement to [19, Section 2.1]. We first introduce the relevant spaces and the de Rham complex in the infinite-dimensional setting and then introduce their finite-dimensional counterparts. For an extensive introduction to function spaces, see [20] or any other textbook on functional analysis.
2.1 The infinite-dimensional setting
@@
They form the well-known de Rham Hilbert complex [21, 22],
$\mathbb{R}\hookrightarrow H^{1}\left(\Omega\right)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}H(\mathrm{curl};\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}H(\mathrm{div};\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}L^{2}(\Omega)\to 0.$
@@
Note that, at the continuous level, $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ are different, see (8), and trace spaces $\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right)$ and $\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right)$ are in a duality, see [23].
2.2 The finite-dimensional setting
Assume that we are given finite-dimensional spaces $G(\Omega)\subset H^{1}(\Omega)$, $C(\Omega)\subset H(\mathrm{curl};\Omega)$, $D(\Omega)\subset H(\mathrm{div};\Omega)$, $S(\Omega)\subset L^{2}(\Omega)$ which form a discretization of the continuous de Rham Hilbert complex,
(9) $\mathbb{R}\hookrightarrow G(\Omega)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}C(\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}D(\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}S(\Omega)\to 0,$
which is also exact in the contractible domain $Omega$. We will also consider the following subspaces with boundary conditions, where $\Gamma\subset\partial\Omega$,
@@ 2.2 The finite-dimensional setting
Moreover, the finite-dimensional trace spaces will be denoted by $\mathcal{T}G(\Omega,\Gamma)$, $\mathcal{T}C(\Omega,\Gamma)$, $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$, $\mathcal{T}D(\Omega,\Gamma)$. Unlike their continuous counterparts, finite-dimensional trace spaces $\mathcal{T}C(\Omega,\Gamma)$ and $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$ usually are different.
@@
which is skew-symmetric with respect to any two of the three entries. Note that, in the infinite-dimensional setting, the $L^{2}$-boundedness of $\left\langle{\bm{\alpha}\times\bm{\beta}},{\bm{\gamma}}\right\rangle_{\Omega}$ is not implied even if $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are $L^{2}$ vector fields. However, in the finite-dimensional setting where the spaces $C(\Omega)$ and $D(\Omega)$ consist of piecewise-polynomial finite element vector fields, it is in fact $L^{2}$-bounded.
@@ 3 Spatially-discrete formulation and its conservation properties
In this section, we present a spatially discrete weak formulation of (2) and analyze its conservation properties. As a remark, for the sake of neatness, we discuss only the three-dimensional formulation and its temporal discretization in this paper. Obtaining their two-dimensional versions is straightforward.
@@ 3.1 Spatially-discrete formulation
To increase the notational clarity, we omit the part for the temporal domain. The proposed semi-discrete weak formulation of (2) in $\mathbb{R}^{3}$ is written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, seek $\left(\bm{u}_{h},\bm{\omega}_{h},P_{h},\bm{H}_{h}\right)\in D_{\hat{u}}(\Omega,\Gamma_{\hat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\hat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
@@
Note that this assumes that the supplied boundary conditions are exactly representable as members of the chosen finite-dimensional spaces; more generally, one can always project the boundary conditions into the appropriate finite-dimensional trace spaces. See (5) and (6) for the general boundary setting.